# Transferring Features Across Language Models With Model Stitching

### v1

@@ Abstract
In this work, we demonstrate that affine mappings between residual streams of language models is a cheap way to effectively transfer represented features between models. We apply this technique to transfer the weights of Sparse Autoencoders (SAEs) between models of different sizes to compare their representations. We find that small and large models learn highly similar representation spaces, which motivates training expensive components like SAEs on a smaller model and transferring to a larger model at a FLOPs savings. For example, using a small-to-large transferred SAE as initialization can lead to 50% cheaper training runs when training SAEs on larger models. Next, we show that transferred probes and steering vectors can effectively recover ground truth performance. Finally, we dive deeper into feature-level transferability, finding that semantic and structural features transfer noticeably differently while specific classes of functional features have their roles faithfully mapped. Overall, our findings illustrate similarities and differences in the linear representation spaces of small and large models and demonstrate a method for improving the training efficiency of SAEs.
@@ 1 Introduction
A popular perspective to view the internal computation of LLMs is through the lens of features, or functions of the input that serve a downstream purpose like predicting the next token or composing to form model circuits [Olah et al., 2020, Huben et al., 2024]. The Linear Representation Hypothesis (LRH) posits that these features are represented as directions in a high dimensional space [Park et al., 2024, Elhage et al., 2022] and sparse dictionary learning methods [Olshausen and Field, 1996, Faruqui et al., 2015, Serre, 2006] such as Sparse Autoencoders (SAEs) have recently gained popularity for interpreting models through this lens [Huben et al., 2024, Bricken et al., 2023, Gao et al., 2025, Kissane et al., 2024]. A second hypothesis, the Strong Model Universality [Li et al., 2015] predicts that good models learn the same data representations.
@@
Our analysis aims to address the faithfulness of the linear representation hypothesis and findings of universality. If two models represent enough features as directions in similarly organized spaces, we should be able to transfer interpretable feature directions between them. In the present work, we propose the use of model stitching as a mapping for learned representations in language models. Our contributions (Figure 1) can be summarized as follows:
@@
    We introduce a compute-efficient method to transfer SAEs (§3) trained on one model to another, which can be used to initialize SAE training runs on larger models with those trained on smaller ones (§3.1), allowing for an 50% FLOPs savings. We also demonstrate the stitch can also be used to transfer probes and steering vectors (§4).
@@
    We apply this method to within-family model pairs in Pythia [Biderman et al., 2023], GPT2 [Radford et al., 2019], and Gemma-2 [Team et al., 2024] and find strong similarity in downstream metrics, supporting universality (§2).
@@
    We perform a fine-grained analysis on feature transfer and analyze how well semantic and structural features transfer using the stitch (§5.1). We also find SAE features representing well-known universal features and note their transferability (§5.2).
@@ 2 Language Model Stitching
Figure 1: Overview of the main methodologies. (a) We train two affine mappings $\mathcal{T}_{\uparrow,\downarrow}$ concurrently to map between the residual streams of two language models $A$ and $B$. The mappings $\mathcal{T}_{\uparrow,\downarrow}$ are then used to transfer (b) the weights of entire SAEs from $A$ to $B$, which (c) give better initializations that save compute when training SAEs on $B$. The approximate “scaling law” for SAE training is shifted to the left when training from the transferred initialization, capturing the intuition that the transferred initialization saves the work of relearning shared features. More generally, the stitches can be used to transfer (d) arbitrary vectors between the residual stream spaces, including probes and steering vectors.
@@
We would like to learn a mapping from the residual stream of language model $A$ to model $B$ at some layers, such that computation can start in the layers of $A$ and finish in $B$ (see Figure 1a). As notation, let $h_{\textbf{LM}}^{\ell}(t)\in\mathbb{R}^{d_{\textbf{LM}}}$ denote the representation in model LM of input token(s) $t$ which has hidden state dimension $d_{\textbf{LM}}$ at layer $\ell$. Model stitching learns a mapping $\mathcal{T}$ (the stitch) between two model-layer pairs $(A,\ell_{A})$ and $(B,\ell_{B})$. $\mathcal{T}$ maps from $h_{A}^{\ell_{A}}$ to $h_{B}^{\ell_{B}}$ informally, we would like $\left[\mathcal{T}\circ h_{A}^{\ell_{A}}\right](t)\approx h_{B}^{\ell_{B}}(t)$ where $\circ$ is function composition. The existence of such a mapping $\mathcal{T}$ relies on universality, especially if $\mathcal{T}$ preserves the hypothesis classes of model $A$ and $B$. We are interested in the particular case when $A$ and $B$ are decoder-only language models from the same family (trained on the same data) but $d_{B}\geq d_{A}$. In this setup (visualized in Figure 1a), we will consider the two stitching mappings $\mathcal{T}_{\uparrow}$, which maps “up” from $A$ to $B$, and $\mathcal{T}_{\downarrow}$, which maps “down” from $B$ to $A$. Furthermore, we will assume that both $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ are affine transformations:
$\mathcal{T}_{\uparrow}:\mathbb{R}^{d_{A}}\to\mathbb{R}^{d_{B}}$ $,\quad\ell_{A}\mapsto\ell_{A}P_{\uparrow}+b_{\uparrow},$ (1)
$\mathcal{T}_{\downarrow}:\mathbb{R}^{d_{B}}\to\mathbb{R}^{d_{A}}$ $,\quad\ell_{B}\mapsto\ell_{B}P_{\downarrow}+b_{\downarrow}.$ (2)
@@
Ideally, we would like $\mathcal{T}_{\{\uparrow,\downarrow\}}$ to be faithful to the downstream objective (i.e., language modeling). In practice, training directly on the next-token prediction objective would involve backpropagating gradients over the back halves of $A$ and $B$, which is unnecessarily expensive. We find that the reconstruction mean squared error (MSE) is a sufficient training objective align these models. Despite the dimensionality gap, we also still desire an almost-invertible transformation i.e. $\mathcal{T}_{\uparrow}\circ\mathcal{T}_{\downarrow}$ should be close to an identity and vice versa. Therefore, we also introduce a regularization penalty that encourages $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ to invert each other with relative strength $\alpha$, finding that the penalty improves fidelity (ablation in §A.2) and has further motivation when we consider transferring SAEs in §3.1 and §B.1. Formally, we train $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ concurrently via the loss function
$\mathcal{L}(t)$ $=\text{MSE}\left(\left[\mathcal{T}_{\uparrow}\circ h^{\ell_{A}}_{A}\right](t),h_{B}^{\ell_{B}}(t)\right)+\text{MSE}\left(\left[\mathcal{T}_{\downarrow}\circ h_{B}^{\ell_{B}}\right](t),h_{A}^{\ell_{A}}(t)\right)$
@@
$+\alpha\text{MSE}\left(\left[\mathcal{T}_{\downarrow}\circ\mathcal{T}_{\uparrow}\circ h_{A}^{\ell_{A}}\right](t),h_{A}^{\ell_{A}}(t)\right)+\alpha\text{MSE}\left(\left[\mathcal{T}_{\uparrow}\circ\mathcal{T}_{\downarrow}\circ h_{B}^{\ell_{B}}\right](t),h_{B}^{\ell_{B}}(t)\right).$ (3)
@@
Table 1: Model-layer pairs that we stitch between and the associated downstream next token CE losses. All activations are taken before the layer’s computation. We primarily stitch from middle residual layers but for specific use cases (e.g., steering) we stitch between $\nicefrac{{3}}{{4}}$ of the depth. The losses are computed stitching from $A$ to $B$, $B$ to $A$, $A$ to $B$ back to $A$, and $B$ to $A$ back to $B$. We calculate relative losses to the model that comes last and color code with blue and purple for model $A$ and model $B$ respectively.
@@ 3 Transferring SAEs
Sparse autoencoders (SAEs) have recently resurfaced in interpretability as a method of decomposing residual stream representations $x\in\mathbb{R}^{d}$ from language models into sparse nonnegative feature activations $f(x;\theta)\in\mathbb{R}^{M}$, with $M\gg d$ and $\theta=(W_{e},b_{e},W_{d},b_{d})$ of the encoder matrix, encoder bias, decoder matrix, and decoder bias respectively that specify the affine transformations in the forward pass, given as
@@
The columns of the encoder matrix are the encoding/detection directions of the features. The rows of the decoder matrix are the decoding/representation directions of the features that the SAE decomposes into. Feature $i$ activates on token $t$ if $f_{i}\left(h^{\ell}(t);\theta\right)>0$. SAEs are trained to reconstruct $x$ while having sparse activations via an objective function consisting of MSE between $x$ and $\textbf{SAE}(x;\theta)$ and, depending on the architecture, an $L_{p}$ regularization (usually with $p=1$) on $f(x;\theta)$ with strength controlled via a parameter $\lambda>0$, respectively. We use pretrained SAEs with TopK activation [Gao et al., 2025] for Pythia [^1] https://github.com/EleutherAI/sparsify and JumpReLU SAEs for Gemma [Lieberum et al., 2024] and train TopK SAEs from scratch on the GPT2 models using SAELens [Bloom et al., 2024]. [^2] Gao et al. [2025] releases pretrained TopK SAEs for GPT2. However, since these SAEs use a LayerNorm normalization scheme but our stitches are trained on unnormalized residual stream activations, the SAE transfer computation does not extend as easily.
@@ 3.1 Efficient SAE Transfer
Equipped with the trained stitch transformations, we make a key observation that an SAE with parameters $\theta=(W_{e},b_{e},W_{d},b_{d})$ trained on layer $\ell_{A}$ in $A$ can be transferred to layer $\ell_{B}$ in $B$. This procedure is illustrated in Figure 1b - specifically, for a latent $h_{B}^{\ell_{B}}(t)$ in model $B$, we (1) transfer down using $T_{\downarrow}$, (2) apply the original SAE parameterized by $\theta$, and (3) transfer back up using $T_{\uparrow}$. Because all the transformations are affine, this forward computation exactly specifies a new SAE on model $B$ parameterized by
@@
We report evaluations of the transferred SAEs in §C, which are better than random but certainly worse than a fully trained SAE on $B$ (e.g., a Fraction Unexplained Variance [FUV] of 0.21 to 0.42 in Gemma 2B to 9B). The metrics simply serve as further evidence of the weak universality in the midst of a representational gap between $A$ and $B$. We note that the weight matrices of the transferred SAE are still rank $\leq d_{A}<d_{B}$ i.e., the SAE is only detecting and writing to features along a rank $d_{A}$ subspace embedded within the rank $d_{B}$ ambient embedding space of model $B$, contributing to their lackluster performance.
@@
After obtaining a transferred SAE on $B$, we investigate how much continued training is required to match the performance of an SAE fully trained on model $B$. This question leads to a key application of the stitch and implicitly model universality: the transferred SAE can be used to initialize a new SAE in model $B$, “saving” the compute of having to relearn features that transfer well. Concretely, consider the problem of training one SAE for two models of different sizes in the same family [Lieberum et al., 2024]. The straightforward approach is to separately train a SAE from scratch on each model. However, we find that given a fully trained SAE on the small model, training a stitch and using the transferred small model SAE in the initialization scheme of the larger SAE allows for training of a comparable SAE to training from scratch but in a cheaper total FLoating point OPerations (FLOPs) budget. We support this claim by training TopK SAEs on the Pythia model pair and repeat with GPT2 in §D.1. We estimate FLOP counts based on the assumption that activations are cached prior to training both the stitch and the SAEs. All SAEs are trained using SAELens [Bloom et al., 2024] and other training and FLOP estimation details are included in §D.
@@ 4 Downstream Applications
In order to probe model $B$, we can transfer the SAE from $A$ to $B$ and use the activations from the same $k$ features in $\mathcal{F}_{k}$ to probe $B$. This procedure is equivalent to computing $\mathcal{T}_{\downarrow}(h_{B}^{\ell_{B}}(t))$, then applying a probe on $\mathcal{F}_{k}$’s activations on the stitched residual stream. Notably, the probe we use can either be completely reused as a pure evaluation of the stitch (i.e. the same probe that we trained when probing model $A$) or retrained to extract the full probing capability of the transferred features.
@@
In this section, we explore transferring steering vectors in a simple task: altering the response language from en to a target language $L$ in EuroParl [Koehn, 2005]. We compute a steering vector over the first $100$ paired examples in each en-$L$ dataset by unit-normalizing the difference in mean activation on $L$ tokens vs. English tokens [Panickssery et al., 2023]. To steer, we clamp the component in the feature’s direction to $\bar{z}$, where we set $\bar{z}$ to the mean of $\langle h^{\ell}(t),v\rangle$ over the tokens in $L$ (i.e., positive examples of the desired behavior). Explicitly, if $v$ is a steering vector and $\|v\|_{2}=1$, then we steer by modifying the hidden state as
$c$ $=\bar{z}-\langle h^{\ell}(x),v\rangle,$ (8)
$h^{\ell}(x)^{\prime}$ $=h^{\ell}(x)+cv.$ (9)

### v2

@@ Abstract
In this work, we demonstrate that affine mappings between residual streams of language models is a cheap way to effectively transfer represented features between models. We apply this technique to transfer the weights of Sparse Autoencoders (SAEs) between models of different sizes to compare their representations. We find that small and large models learn similar representation spaces, which motivates training expensive components like SAEs on a smaller model and transferring to a larger model at a FLOPs savings. In particular, using a small-to-large transferred SAE as initialization can lead to 50% cheaper training runs when training SAEs on larger models. Next, we show that transferred probes and steering vectors can effectively recover ground truth performance. Finally, we dive deeper into feature-level transferability, finding that semantic and structural features transfer noticeably differently while specific classes of functional features have their roles faithfully mapped. Overall, our findings illustrate similarities and differences in the linear representation spaces of small and large models and demonstrate a method for improving the training efficiency of SAEs.
@@ 1 Introduction
A popular perspective to view the internal computation of LLMs is through the lens of features, or functions of the input that serve a downstream purpose like predicting the next token or composing to form model circuits [Olah et al., 2020, Huben et al., 2024]. The Linear Representation Hypothesis (LRH) posits that these features are represented as directions in a high dimensional space [Park et al., 2024, Elhage et al., 2022] and sparse dictionary learning methods [Olshausen and Field, 1996, Faruqui et al., 2015, Serre, 2006] such as Sparse Autoencoders (SAEs) have recently gained popularity for interpreting models through this lens [Huben et al., 2024, Bricken et al., 2023, Gao et al., 2025, Kissane et al., 2024]. A second hypothesis, Strong Model Universality [Li et al., 2015], predicts that good models learn the same data representations.
@@
Our work relies on the ansatz that if two models represent enough features as directions in similarly organized spaces, we should be able to transfer interpretable linear features between them. We propose the use of model stitches as the mappings for learned representations between language models. Our contributions (Figure 1) can be summarized as follows:
@@
    We learn stitches to transfer SAEs (§3) trained on one model to another, which are used to initialize SAE training runs on larger models with those trained on smaller ones (§3.1), allowing for an 50% FLOPs savings. We also demonstrate that stitches can be used to transfer probes and steering vectors (§4).
@@
    The stitches are affine mappings trained on within-family model pairs in Pythia [Biderman et al., 2023], GPT2 [Radford et al., 2019], and Gemma-2 [Team et al., 2024]. We find downstream metrics are preserved by the mappings, supporting weak universality (§2).
@@
    We perform a fine-grained analysis on feature transfer and analyze how well semantic and structural features transfer using the stitch (§5.1). We also find SAE features representing well-known universal functional features and note their transferability (§5.2).
@@ 2 Language Model Stitching
Figure 1: Overview of the main methodologies. (a) We train two affine mappings $\mathcal{T}_{\uparrow,\downarrow}$ concurrently to map between the residual streams of two language models $A$ and $B$. The mappings $\mathcal{T}_{\uparrow,\downarrow}$ are then used to transfer (b) the weights of entire SAEs from $A$ to $B$, which (c) give better initializations that save compute when training SAEs on $B$. The approximate “scaling law” for SAE training is shifted to the left when training from the transferred initialization, capturing the intuition that the transferred initialization saves the work of relearning shared features. More generally, the stitches can be used to transfer (d) arbitrary vectors (probes, steering vectors) between the residual stream spaces.
@@
We would like to learn a mapping from the residual stream of language model $A$ to model $B$ at some layers, such that computation can start in the layers of $A$ and finish in $B$ (see Figure 1a). As notation, let $\textbf{LM}^{\ell}(t)\in\mathbb{R}^{d_{\textbf{LM}}}$ denote the representation in model LM of input token(s) $t$ which has hidden state dimension $d_{\textbf{LM}}$ at layer $\ell$. Model stitching learns a mapping $\mathcal{T}$ (the stitch) between two model-layer pairs $(A,\ell_{A})$ and $(B,\ell_{B})$. Informally, we would like $\left[\mathcal{T}\circ A^{\ell_{A}}\right](t)\approx B^{\ell_{B}}(t)$ where $\circ$ is function composition. The existence of such a mapping $\mathcal{T}$ relies on at least weak universality, especially if $\mathcal{T}$ preserves the hypothesis classes of model $A$ and $B$. We are interested in the particular case when $A$ and $B$ are decoder-only language models from the same family (trained on the same data) but $d_{B}\geq d_{A}$. In this setup (visualized in Figure 1a), we will consider the two stitching mappings $\mathcal{T}_{\uparrow}$, which maps “up” from $A$ to $B$, and $\mathcal{T}_{\downarrow}$, which maps “down” from $B$ to $A$. Furthermore, we will assume that both $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ are affine transformations:
$\mathcal{T}_{\uparrow}:\mathbb{R}^{d_{A}}\to\mathbb{R}^{d_{B}}$ $,\quad h_{A}\mapsto h_{A}P_{\uparrow}+b_{\uparrow},$ (1)
$\mathcal{T}_{\downarrow}:\mathbb{R}^{d_{B}}\to\mathbb{R}^{d_{A}}$ $,\quad h_{B}\mapsto h_{B}P_{\downarrow}+b_{\downarrow},$ (2)
@@
where $h_{A}$ and $h_{B}$ are activations from $A$ and $B$, $P_{\uparrow}\in\mathbb{R}^{d_{A}\times d_{B}}$, $b_{\uparrow}\in\mathbb{R}^{d_{B}}$, $P_{\downarrow}\in\mathbb{R}^{d_{B}\times d_{A}}$, and $b_{\downarrow}\in\mathbb{R}^{d_{A}}$. Ideally, we would like $\mathcal{T}_{\{\uparrow,\downarrow\}}$ to be faithful to the downstream objective (i.e., language modeling). In practice, training directly on the next-token prediction objective would involve backpropagating gradients over the back halves of $A$ and $B$, which is unnecessarily expensive. We find that the reconstruction mean squared error (MSE) is a sufficient training objective to align the models. Despite the dimensionality gap, we also still desire “almost”-invertible transformations i.e. $\mathcal{T}_{\uparrow}\circ\mathcal{T}_{\downarrow}$ should be close to an identity and vice versa. Therefore, we also introduce a regularization penalty that encourages $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ to invert each other with relative strength $\alpha$, finding that the penalty improves fidelity (ablation in §A.2) and has further motivation when we consider transferring SAEs in §3.1 and §B.1. Formally, we train $\mathcal{T}_{\uparrow}$ and $\mathcal{T}_{\downarrow}$ concurrently via the loss function on a token $t$
$\mathcal{L}(t)$ $=\text{MSE}\left(\left[\mathcal{T}_{\uparrow}\circ A^{\ell_{A}}\right](t),B^{\ell_{B}}(t)\right)+\text{MSE}\left(\left[\mathcal{T}_{\downarrow}\circ B^{\ell_{B}}\right](t),A^{\ell_{A}}(t)\right)$
@@
$+\alpha\text{MSE}\left(\left[\mathcal{T}_{\downarrow}\circ\mathcal{T}_{\uparrow}\circ A^{\ell_{A}}\right](t),A^{\ell_{A}}(t)\right)+\alpha\text{MSE}\left(\left[\mathcal{T}_{\uparrow}\circ\mathcal{T}_{\downarrow}\circ B^{\ell_{B}}\right](t),B^{\ell_{B}}(t)\right).$ (3)
@@
Table 1: Model-layer pairs that we stitch between and the associated downstream next token CE losses on OpenWebText. All activations are taken before the layer’s computation. We primarily stitch from middle residual layers but for specific use cases (e.g., steering) we stitch between $\nicefrac{{3}}{{4}}$ of the depth. The losses are computed stitching from $A$ to $B$, $B$ to $A$, $A$ to $B$ back to $A$, and $B$ to $A$ back to $B$. We calculate relative losses to the model that comes last and color code with blue and purple for model $A$ and model $B$ respectively.
@@ 3 Transferring SAEs
Sparse autoencoders (SAEs) have recently resurfaced in interpretability as a method of decomposing residual stream representations $x\in\mathbb{R}^{d}$ from a language model LM at layer $\ell$ into sparse nonnegative feature activations $f(x;\theta)\in\mathbb{R}^{M}$, with $M\gg d$ and $\theta=(W_{e},b_{e},W_{d},b_{d})$ of the encoder matrix, encoder bias, decoder matrix, and decoder bias respectively that specify the affine transformations in the forward pass, given as
@@
The columns of the encoder matrix are the encoding/detection directions of the features. The rows of the decoder matrix are the decoding/representation directions of the features that the SAE decomposes into. Feature $i$ activates on token $t$ if $f_{i}\left(\textbf{LM}^{\ell}(t);\theta\right)>0$. SAEs are trained to reconstruct $x$ while having sparse activations via an objective function consisting of MSE between $x$ and $\textbf{SAE}(x;\theta)$ and, depending on the architecture, an $L_{p}$ regularization (usually with $p=1$) on $f(x;\theta)$ with strength controlled via a parameter $\lambda>0$, respectively. We use pretrained SAEs with TopK activation [Gao et al., 2025] for Pythia [^1] https://github.com/EleutherAI/sparsify and JumpReLU SAEs for Gemma [Lieberum et al., 2024] and train TopK SAEs from scratch on the GPT2 models using SAELens [Bloom et al., 2024]. [^2] Gao et al. [2025] releases pretrained TopK SAEs for GPT2. However, since these SAEs use a LayerNorm normalization scheme but our stitches are trained on unnormalized residual stream activations, the SAE transfer computation does not extend as easily.
@@ 3.1 Efficient SAE Transfer
Equipped with the trained stitch transformations, we make a key observation that an SAE with parameters $\theta=(W_{e},b_{e},W_{d},b_{d})$ trained on layer $\ell_{A}$ in $A$ can be transferred to layer $\ell_{B}$ in $B$. This procedure is illustrated in Figure 1b - specifically, for a latent $h_{B}=B^{\ell_{B}}(t)$ in model $B$, we (1) transfer down using $T_{\downarrow}$, (2) apply the original SAE parameterized by $\theta$, and (3) transfer back up using $T_{\uparrow}$. Because all the transformations are affine, this forward computation exactly specifies a new SAE on model $B$ parameterized by
@@
We report evaluations of the transferred SAEs in §C, which are better than random but certainly worse than a fully trained SAE on $B$ (e.g., a Fraction Unexplained Variance [FUV] of 0.21 to 0.42 in Gemma 2B to 9B). The metrics serve as evidence of weak universality in the midst of a representational gap between $A$ and $B$. As a partial explanation, we note that the weight matrices of the transferred SAE are still rank $\leq d_{A}<d_{B}$ i.e., the SAE is only detecting and writing to features along a rank $d_{A}$ subspace embedded within the rank $d_{B}$ ambient embedding space of model $B$, contributing to their lackluster performance.
@@
After obtaining a transferred SAE on $B$, we investigate how much continued training is required to match the performance of an SAE fully trained on model $B$. This question leads to a key application of the stitch and implicitly weak universality: the transferred SAE can be used to initialize a new SAE in model $B$, “saving” the compute of having to relearn features that transfer well. Concretely, consider the problem of training one SAE for two models of different sizes in the same family [Lieberum et al., 2024]. The straightforward approach is to separately train a SAE from scratch on each model. However, we find that given a fully trained SAE on the small model, training a stitch and using the transferred small model SAE in the initialization scheme of the larger SAE allows for training of a comparable SAE to training from scratch but in a cheaper total FLoating point OPerations (FLOPs) budget. We support this claim by training TopK SAEs on the Pythia model pair and repeat with GPT2 in §D.1. We estimate FLOP counts based on the assumption that activations are cached prior to training both the stitch and the SAEs. All SAEs are trained using SAELens [Bloom et al., 2024] and other training and FLOP estimation details are included in §D.
@@ 4 Downstream Applications
In order to probe model $B$, we can transfer the SAE from $A$ to $B$ and use the activations from the same $k$ features in $\mathcal{F}_{k}$ to probe $B$. This procedure is equivalent to computing $\mathcal{T}_{\downarrow}(B^{\ell_{B}}(t))$, then applying a probe on $\mathcal{F}_{k}$’s activations on the stitched residual stream. Notably, the probe we use can either be completely reused as a pure evaluation of the stitch (i.e. the same probe that we trained when probing model $A$) or retrained to extract the full probing capability of the transferred features.
@@
In this section, we explore transferring steering vectors in a simple task: altering the response language from en to a target language $L$ in EuroParl [Koehn, 2005]. We compute a steering vector over the first $100$ paired examples in each en-$L$ dataset by unit-normalizing the difference in mean activation on $L$ tokens vs. English tokens [Panickssery et al., 2023]. To steer, we clamp the component in the feature’s direction to $\bar{z}$, where we set $\bar{z}$ to the mean of $\langle\textbf{LM}^{\ell}(t),v\rangle$ over the tokens in $L$ (i.e., positive examples of the desired behavior). Explicitly, if $v$ is a steering vector and $\|v\|_{2}=1$, then we steer by modifying the hidden state at a token $t$ as
$c$ $=\bar{z}-\langle\textbf{LM}^{\ell}(t),v\rangle,$ (8)
$\textbf{LM}^{\ell}(t)^{\prime}$ $=\textbf{LM}^{\ell}(t)+cv.$ (9)